You will get $450,000 from your rich uncle Jim in 15 years. The discount rate and the interest rate are 2%/year, compounded annually. What is the present value of this gift from your uncle.

1. $330,356.63

2. $331,356.63

3. $332,256.63

4. $333,256.63

5. $334,356.63

The retirement benefit program from your employer will send you 240 monthly checks of $5,000 after you retire. The first check will be a month from your retirement. The interest and discount rates are 0.2%/month, compounded monthly. If you retire today, then what is the present value of this benefit program?

1. $925,299.80

2. $940,299.80

3. $949,299.80

4. $952,299.80

5. $955,299.80
Answer #1

The solution to the given question is :

1.
Option 5)
=450000/1.02^15=334356.6285


2.
Option 4)
=5000/0.2%*(1-1/1.002^240)=952299.7912